OpenStudy (anonymous):

At what point does the normal to y=0−1x−1x2 at (1,−2) intersect the parabola a second time?

OpenStudy (tkhunny):

Have you considered writing the equation of the Normal Line?

OpenStudy (anonymous):

??

OpenStudy (tkhunny):

You cannot solve this problem if you do not know what Normal Line is. It is related to the Tangent Line. Can you find its equation? We have a point, (1,-2). All we need is a slope.

OpenStudy (anonymous):

The normal line is perpendicular to the tangent line. If two lines are perpendicular their slopes are negative reciprocals -- i.e. if the slope of the first line is m then the slope of the second line is −1/m

OpenStudy (tkhunny):

Okay, what is the slope of the tangent lines at (1,-2)?

OpenStudy (anonymous):

-1

OpenStudy (tkhunny):

y= −x − x^2 at (1,−2) y' = -1 - 2x y'(1) = -1 - 2(1) = -3 -- This is the slope of the tangent line through (1,-2). -3(1/3) = -1 1/3 is the slope of the Normal Line through (1,-2) Where did you get -1? You didn't say what that was supposed to be, so it could have been anything. Not really clear where it came from.